OpenStudy (anonymous):

Use the limit definition to compute f'(a): f(x) = 1x + x−1, a = 4 Find an equation of the tangent line using variables y and x.

OpenStudy (anonymous):

Sorry! The equation is f(x) = 1x + x^(−1)

OpenStudy (anonymous):

aka \[x+\frac{1}{x}\]?

OpenStudy (anonymous):

f(x) = 1x + x^−1, a = 4 f '(x)=1+x^-2 f '(4)=1+4^-2=17/16=1.0625

OpenStudy (anonymous):

Satellite: Yes. It's 15/16 actually. I almost understand why, but I'm still having some trouble with the algebra.

OpenStudy (precal):

Use the power rule. \[f(x)=x+x^-1, a=4\] \[f \prime(x)=1-1x ^{-2}\] \[f\prime(4)=1-(1/4^2)\]=(15/16)